OpenStudy (anonymous):

24. The perimeter of a rectangular concrete slab is 114 feet and its area is 702 square feet. Find the dimensions of the rectangle. a. Using l for the length of the rectangle, write an expression for the width of the rectangle in terms of l. (Hint: Solve the formula for w.) Show your work. b. Write a quadratic equation using l, the expression you found in part (a), and the area of the slab. c. Solve the quadratic equation. Use the two solutions to find the dimensions of the rectangle

OpenStudy (campbell_st):

well 1st you need to halve the perimeter to find the dimensions... the reason is  so 54 = l + w the area is \[702 = l \times w\] so you will need to solve the 2 equations simultaneously using 54 = 1 + w make w the subject w = 54 - l and subsitute into the area formula \[702 = l \times(54 - l)\] when you distribute, you'll find you need to solve a quadratic equation.. hope this helps.

OpenStudy (anonymous):

thank you so much